# Image-Text-Image Knowledge Transferring for Lifelong Person Re-Identification with Hybrid Clothing States

### v1

@@ 3.1. Formulation and Overview
Task definition. In the task of lifelong person ReID with hybrid clothing states, a stream of datasets $\mathcal{D}={\{\mathcal{D}^{(t)}\}}^{T}_{t=1}$ are sequentially collected in different domains. $\mathcal{D}^{(t)}$ denotes the $t$-th added ReID dataset, which could be either the Same-Cloth scenario (i.e., each pedestrian always wears the same clothes in different camera views), or the Cloth-Changing setting (i.e., images of the same person can be captured appearing different clothing). It is a more practical yet challenging task than previous LReID (Ge et al., 2022; Pu et al., 2021; Yu et al., 2023), since the model is required to perform well on seen or unseen domains with various clothing conditions, pursuing the ability to not only defend the catastrophic forgetting problem, but also balance the learned cloth-related and cloth-irrelevant identity knowledge. At the $t$-th step, $\mathcal{D}^{(t)}=\{\mathcal{D}^{(t)}_{train},\mathcal{D}^{(t)}_{test}\}$ contains the training set and the testing set, each of which includes a variety of pairs of images $x_{i}^{(t)}$ and identity labels $y_{i}^{(t)}$. An image encoder $\mathcal{G}_{t-1}$ from the $(t-1)$-th step is thus trained on the training set, to incrementally acquire new knowledge on the $t$-th dataset, that is, $\mathcal{G}_{t-1}\to\mathcal{G}_{t}$. Notably, samples of the previous datasets are not available at the current step due to data privacy and storage limitations. Eventually, the ReID model $\mathcal{G}_{T}$ should perform well on all seen domains, which is evaluated on the testing sets $\{\mathcal{D}^{(t)}_{test}\}^{T}_{t=1}$ at the end of the lifelong learning process of totally $T$ steps. Meanwhile, $\mathcal{G}_{T}$ should generalize well on the testing sets of unseen domains $\{\mathcal{D}^{(u)}_{u\_test}\}^{U}_{u=1}$, where $U$ is the total number of unseen datasets with various clothing states.
@@
Method overview. We propose a new framework Teata, to effectively align, transfer and accumulate knowledge using the text space as an intermediate medium. The overview of our framework is illustrated in Fig. 2. Inspired by CLIP-ReID (Li et al., 2023), we adopt CLIP (Radford et al., 2021) to bridge image and text latent space, and build a closed loop of “image-text-image”. Teata starts by aligning image features with text embeddings to transfer visual information and category labels to the text space. Next, we leverage the robustness and consistency of text semantics to guide the discriminative feature learning of visual representations. This text-to-image knowledge transfer is finally reflected in the parameter update of the image encoder.
@@
Based on these two alternating stages, knowledge is continuously transferred in the image and text space, thereby achieving the accumulation of knowledge. To achieve effective knowledge transfer, we further design a Structured Semantic Prompt (SSP) learning in the first stage, which decomposes the original text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description (see Sec. 3.2). To further alleviate the knowledge confusion caused by incremental learning from same-cloth ReID and cloth-changing ReID tasks, we introduce a Knowledge Adaptation and Projection strategy (KAP) in the second stage, which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting (see Sec. 3.3). Lastly, we describe the details of training and testing procedures of Teata in Sec. 3.4.
3.2. Structured Semantic Prompt Learning
@@
CLIP-ReID (Li et al., 2023) utilizes a series of learnable tokens ${[X]}_{1}{[X]}_{2}\dots[X]_{M}$ as a container to distill visual knowledge from images. It is a free-from and task-driven learning process, as the features of each token learned could be different depending on different tasks or datasets. Although this token design is flexible, it could have a negative impact when performing lifelong learning across ReID datasets with different clothing states. Knowledge acquired from same-cloth datasets tends to focus on the visual appearance features, while it requires more fine-grained discriminative clues, such as shoe style, gender and body shape, when applying ReID models to the cloth-changing scenario. This mismatch, in terms of knowledge granularity, would cause a radically different update direction for each token, resulting in an inability to accumulate and transfer knowledge effectively.
@@
Given a person image, we can describe its identity with text, for example, “a middle age man with short hair wearing a white skirt and black pants”. We observe that the description of attributes has a similar form of [content object]. Moreover, for instance, [short hair] is fine-grained information helpful for both same-cloth and cloth-changing ReID, and [white skirt] is appearance information that is more discriminative under the condition of changing clothes. It thus motivates us to present a Structured Semantic Prompt (SSP) with shared and specific tokens. Shared tokens are expected to function as [object] and be shared across all tasks so that the text description has similar granularity. The specific tokens represent [content] information unique to each identity, tailored for different ReID tasks. It is an ideal intuition, but unfortunately limited by the absence of attribute labels. To this end, we build upon learnable tokens and further structurally decompose them in the text semantics. More specifically, we assign the structured text prompt $P_{j}^{(t)}$, $j\in[1,N^{(t)}]$, to each pedestrian as follows.
(1) $P_{j}^{(t)}={\rm A\ photo\ of\ a}\ [{X}_{1}\ {Y}_{1}]\cdots[{X}_{M}\ {Y}_{M}]\ {\rm person.}$
@@
where $M$ means the total number of token pairs, ${X}_{m}$ and ${Y}_{m}$ $(m\in[1,M])$ denote the specific and shared tokens, respectively. They are all learnable parameters with the same dimension and are alternately combined to form each structured pair $[{X}_{m}\ {Y}_{m}]$.
@@
With our proposed structured semantic prompt, models are able to integrate knowledge across domains with various clothing states in the unified textual semantic space. Analyses in Sec. 4.4 and Tab. 6 further support its efficacy, albeit in a straightforward form. To optimize it, we adopt CLIP contrastive losses $\mathcal{L}_{i2t}$ and $\mathcal{L}_{t2i}$ (Radford et al., 2021) which constrain the bidirectional alignment of image features and text embeddings:
(2) $\mathcal{L}_{stage1}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}$
3.3. Knowledge Adaptation and Projection
@@
Once the first stage of structured semantic prompt learning is done, we expect to transfer the knowledge from text to image space, so as to update the parameters of the image encoder $\mathcal{G}_{t}$. However, as shown in Fig. 2, it is noteworthy that we use the previous model of $\mathcal{G}_{t-1}$ to extract features of samples from the $t$-th dataset, and then distill knowledge to text embeddings. In the LReID-Hybrid task, the knowledge acquired from ReID tasks with different clothing states may be contradictory. For example, the visual appearance features are discriminate representations for same-cloth ReID, but may be misleading for cloth-changing ReID. This inevitably leads to learned text embeddings being sub-optimal or not applicable, thus affecting effective knowledge transfer and accumulation. To this end, we improve the second stage of learning with the Knowledge Adaptation and Projection (KAP) strategy.
@@
Adaptation with slow-paced learner. After the first stage, text embeddings for each identity can be obtained using the learned structured semantic prompts and the frozen text encoder. Formally,
(3) $T^{(t)}={\left\{{f^{T}_{j}}^{(t)}=\mathcal{T}(P_{j}^{(t)})\ \big{|}\ j\in[1,N^{(t)}]\right\}}$
@@
As discussed above, instead of freezing text embeddings $T^{(t)}$, we propose to fine-tune them with a slow-paced learner. It shows two benefits. First, we can continuously update text embeddings to make them more suitable for the current ReID task. Meanwhile, the slow-paced learner can prevent the learning of text embeddings from overfitting to the current task, resulting in forgetting previous knowledge. Note that when updating text embeddings, the text encoder $\mathcal{T}$ is kept frozen to reserve the learned text semantic knowledge and the structured semantic prompts are tuned for knowledge adaptation.
@@
Project-then-learn. To transfer the knowledge from text to image space effectively, we use the text embeddings to guide the learning of $\mathcal{G}_{t}$ by projecting image features on the text space and then updating the ReID model, i.e., the image encoder, to accumulate new knowledge. Formally,
(4) $\mathcal{L}_{id}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp(f^{(t)}_{i}\cdot{{f^{S}_{j}}^{(t)}}^{\top})}{\sum\limits_{k=1}^{N^{(t)}}\exp(f^{(t)}_{i}\cdot{{f^{S}_{k}}^{(t)}}^{\top})}$
(5) $q_{j}=\left\{\begin{aligned} &1-\epsilon+\frac{\epsilon}{N^{(t)}}&&,j=y_{i}^{(t)}\\ &\frac{\epsilon}{N^{(t)}}&&,{\rm otherwise}\end{aligned}\right.$
@@
where $f_{i}^{(t)}$ means the image feature of the $i$-th image with the identity label $y_{i}^{(t)}$, and ${f^{S}}^{(t)}$ is the adapted text embeddings and initialized with the L2 normalized text embeddings $T^{(t)}$, and ${f_{j}^{S}}^{(t)}$ denotes the adapted text embedding of the $j$-th identity. $N^{(t)}$ denotes the number of identities in the $t$-th dataset. $\epsilon$ is a small constant for label smoothing regularization (Szegedy et al., 2015), which is simply set to 0.1. Notably, Eq. 4 is effective in fine-tuning the image encoder $\mathcal{G}_{t}$ together with text embeddings. As a result, with the help of the slow-paced learner, both image encoder and text embeddings can be slowly updated so that new knowledge is accumulated without forgetting the old. We still keep the original projection loss $\mathcal{L}_{proj}$ proposed in (Li et al., 2023) to strengthen the learning of the image encoder, because it experimentally brings performance gains.
(6) $\mathcal{L}_{proj}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp(f^{(t)}_{i}\cdot{{f^{T}_{j}}^{(t)}}^{\top})}{\sum\limits_{k=1}^{N^{(t)}}\exp(f^{(t)}_{i}\cdot{{f^{T}_{k}}^{(t)}}^{\top})}$
@@
Alternative variant. Benefiting from CLIP and our proposed SSP, text and image semantics are aligned in the shared latent space. Intuitively, image representations that are well aligned with text embeddings can also be used in Eq. 4 to guide knowledge transfer and accumulation. Specifically, we can average all the extracted image features of each identity to obtain image representations $V^{(t)}$ at the beginning of the second stage training. Formally,
(7) $V^{(t)}={\left\{{f^{V}_{j}}^{(t)}=\frac{1}{\left|X^{(t)}_{j}\right|}\sum\limits_{f^{(t)}\in X^{(t)}_{j}}f^{(t)}\ \bigg{|}\ j\in\left[1,N^{(t)}\right]\right\}}$
where $X^{(t)}_{j}$ denotes the set of all person images with the same identity label $j$, and |||*|| ∗ | denotes the cardinal number of the set. Then, L2 normalized $V^{(t)}$ is used for the initialization of the classifier in Eq. 4. In Sec. 4.4, we thoroughly investigate the efficacy of our designs, including knowledge adaptation with text embeddings (w/ KA-T), text-aligned image representations (w/ KA-V), random initialization (by default), and slow-paced learner (w/ SL).
3.4. Details of Training and Inference
Overall, two stages are involved in each step $t$ during the lifelong evolution of the model. In the first stage, both image encoder $\mathcal{G}_{t-1}$ and text encoder $\mathcal{T}$ and frozen, and Eq. 2 is adopted for the optimization of SSP. In the second stage, $\mathcal{G}_{t}$ is optimized with the guidance of the learned text embeddings. The losses involved in the second stage are formulated as follows.
(8) $\mathcal{L}_{stage2}=\lambda_{1}\mathcal{L}_{proj}+\lambda_{2}(\mathcal{L}^{b}_{id}+\mathcal{L}_{id})+\lambda_{3}(\mathcal{L}^{b}_{tri}+\mathcal{L}_{tri})$
wherein $\mathcal{L}^{b}_{id}$ and $\mathcal{L}^{b}_{tri}$ denote the identity classification loss (Luo et al., 2019) and triplet loss (Hermans et al., 2017) applied on the image features before the last projecting layer of the CLIP image encoder. $\mathcal{L}_{id}$ and $\mathcal{L}_{tri}$ are applied to the representations after projection. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are coefficients to balance between different losses. After training in step $t$, only the image encoder $\mathcal{G}_{t}$ is used for ReID evaluation. The cosine distances between two person images as measured during inference.
4. Experiments
4.1. Datasets and Evaluation Protocol
Seen datasets. Two widely-used conventional same-cloth datasets, i.e., Market-1501 (Zheng et al., 2015) and MSMT17 (Wei et al., 2018), and two widely-used cloth-changing datasets, i.e., LTCC (Qian et al., 2020) and PRCC (Yang et al., 2019), are used in our proposed LReID-Hybrid benchmarks for training. The dataset statistics are shown in Tab. 1. We design six training orders to imitate different situations comprehensively as follows.

### v2

@@ III-A1 Task Definition
In the task of lifelong person ReID with hybrid clothing states, a stream of datasets $\mathcal{D}={\{\mathcal{D}^{(t)}\}}^{T}_{t=1}$, where $T$ is the total number of datasets, are sequentially collected in different domains. $\mathcal{D}^{(t)}$ denotes the $t$-th added ReID dataset, which could be either the Same-Cloth scenario (i.e., each pedestrian always wears the same clothes in different camera views), or the Cloth-Changing setting (i.e., images of the same person can be captured appearing different clothing). It is a more practical yet challenging task than previous LReID [2, 1, 3], since the model is required to perform well on seen or unseen domains with various clothing conditions, pursuing the ability to not only solve the catastrophic forgetting problem, but also balance the learned cloth-related and cloth-irrelevant identity knowledge. At the $t$-th step, $\mathcal{D}^{(t)}=\{\mathcal{D}^{(t)}_{train},\mathcal{D}^{(t)}_{test}\}$ contains the training set and the testing set, each of which includes a variety of pairs of image $x_{i}^{(t)}$ and identity label $y_{i}^{(t)}$. An image encoder $\mathcal{G}_{t-1}$ from the (t1)(t-1)-th step is thus trained on the training set, to incrementally acquire new knowledge on the $t$-th dataset, that is, $\mathcal{G}_{t-1}\to\mathcal{G}_{t}$. Notably, samples of the previous datasets are not available at the current step due to data privacy and storage limitations. Eventually, the ReID model $\mathcal{G}_{T}$ should perform well on all seen domains, which is evaluated on the testing sets $\{\mathcal{D}^{(t)}_{test}\}^{T}_{t=1}$ at the end of the lifelong learning process of totally $T$ steps. Meanwhile, $\mathcal{G}_{T}$ should generalize well on the testing sets of unseen domains $\{\mathcal{D}^{(u)}_{u\_test}\}^{U}_{u=1}$, where $U$ is the total number of unseen datasets with various clothing states.
III-A2 Method Overview
@@
We propose a new framework Teata, to effectively align, transfer, and accumulate knowledge using the text space as an intermediate medium. The overview of Teata is illustrated in Fig. 2. Inspired by CLIP-ReID [12], we adopt CLIP [11] to bridge the image and text latent space, and build a closed loop of “image-text-image”. We start by aligning image features with text embeddings to transfer visual information and category labels to the text space. Next, we leverage the robustness and consistency of text semantics to guide the discriminative feature learning of visual representations. This text-to-image knowledge transfer is finally reflected in the parameter update of the image encoder.
@@
Based on these two alternating stages, knowledge is continuously transferred in the image and text space, thereby achieving the accumulation of knowledge. To achieve effective knowledge transfer, we further design a Structured Semantic Prompt (SSP) learning in the first stage, which decomposes the original text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description (see Sec. III-B). To further alleviate the knowledge confusion caused by incremental learning from same-cloth ReID and cloth-changing ReID tasks, we introduce a Knowledge Adaptation and Projection strategy (KAP) in the second stage, which tunes text knowledge via a slow-paced learner and adapts the image encoder to different tasks without catastrophic forgetting (see Sec. III-C). Lastly, we describe the details of training and testing procedures of Teata in Sec. III-D.
III-B Structured Semantic Prompt Learning
@@
CLIP-ReID [12] utilizes a series of learnable tokens ${[X]}_{1}{[X]}_{2}\dots[X]_{M}$ as a container to distill visual knowledge from images. It is a free-from and task-driven learning process, as the features of each token learned could be different depending on different tasks or datasets. Although this token design is flexible, it could have a negative impact when performing lifelong learning across ReID datasets with different clothing states. Knowledge acquired from same-cloth datasets tends to focus on the visual appearance features, while it requires more fine-grained discriminative clues, such as shoe style, gender, and body shape, when applying ReID models to the cloth-changing scenario. This mismatch, in terms of knowledge granularity, would cause a radically different update direction for each token, resulting in an inability to accumulate and transfer knowledge effectively.
@@
Given a person image, we can describe its identity with text, for example, “a middle age man with short hair wearing a white shirt and black pants”. We observe that the description of attributes has a similar form of [content object]. Moreover, [middle age] and [short hair] are fine-grained identity information helpful for both same-cloth and cloth-changing ReID, while [white shirt] and [black pants] are cloth-related appearance information that is more discriminative under the condition of same-cloth ReID. It thus motivates us to present a Structured Semantic Prompt (SSP) with shared and specific tokens. Shared tokens are expected to function as [object] and be shared across all tasks so that the text description has similar granularity. The specific tokens represent [content] information unique to each identity, tailored for different ReID tasks. It is an ideal intuition, but unfortunately limited by the absence of attribute labels. To this end, we build upon learnable tokens and further structurally decompose them in the text semantics. More specifically, we assign a structured text prompt $P_{j}^{(t)}$ to each pedestrian, where $j\in[1,N^{(t)}]$ and $N^{(t)}$ denotes the number of identities in the $t$-th dataset.
$P_{j}^{(t)}={\rm A\ photo\ of\ a}\ [{X}_{1}\ {Y}_{1}]\cdots[{X}_{M}\ {Y}_{M}]\ {\rm person.},$ (1)
@@
where $M$ means the total number of token pairs, ${X}_{m}$ and ${Y}_{m}$ $(m\in[1,M])$ denote the specific and shared tokens, respectively. They are all learnable parameters with the same dimension and are alternately combined to form each structured pair ${X}_{m}\ {Y}_{m}]$. As discussed above, $\left\{X_{m}\right\}_{m=1}^{M}$ are not only dataset-specific but also identity-specific, that is, different people have different $\left\{X_{m}\right\}_{m=1}^{M}$. On the other hand, $\left\{Y_{m}\right\}_{m=1}^{M}$ are reused for all identities from different datasets.
@@
With our proposed structured semantic prompt, models can integrate knowledge across domains with various clothing states in the unified textual semantic space. Analyses in Sec. IV-D and Tab. XI further support its efficacy, albeit in a straightforward form. To optimize it, we adopt CLIP contrastive losses $\mathcal{L}_{i2t}$ and $\mathcal{L}_{t2i}$ [11, 12] which constrain the bidirectional alignment of image features and text embeddings:
$\mathcal{L}_{stage1}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}.$ (2)
Overall, we use learnable text prompts as containers for image-to-text distillation, and decompose the prompt into several structured pairs. This is an “implicitly feature learning with explicitly structural design” approach to unify knowledge from different domains with the same granularity.
III-C Knowledge Adaptation and Projection
@@
Once the first stage of structured semantic prompt learning is done, we expect to transfer the knowledge from text to image space, so as to update the parameters of the image encoder $\mathcal{G}_{t}$. However, as shown in Fig. 2, the previous model of $\mathcal{G}_{t-1}$ is used to extract features of samples from the $t$-th dataset, and then distill knowledge to text embeddings. In the LReID-Hybrid task, the knowledge acquired from ReID tasks with different clothing states may be contradictory. For example, the visual appearance features are discriminate representations for same-cloth ReID, but may be misleading for cloth-changing ReID. This inevitably leads to learned text embeddings being sub-optimal or not applicable, thus affecting effective knowledge transfer and accumulation. To alleviate the challenge of knowledge representation mismatch, we improve the second stage of learning with the Knowledge Adaptation and Projection (KAP) strategy.
III-C1 Adaptation with Slow-Paced Learner
@@
After the first stage, text embeddings for each identity can be obtained using the learned structured semantic prompts and the frozen text encoder. Formally,
$\mathbf{T}^{(t)}=\Bigg\{{f^{T}_{j}}^{(t)}=\mathcal{T}\big(P_{j}^{(t)}\big)\ \bigg|\ j\in\left[1,N^{(t)}\right]\Bigg\},$ (3)
@@
where $\mathcal{T}$ denotes the text encoder. As discussed above, instead of freezing text embeddings $\mathbf{T}^{(t)}$, we propose to fine-tune them with a slow-paced learner. It shows two benefits. First, we can continuously update text embeddings to make them more suitable for the current ReID task. Meanwhile, the slow-paced learner can prevent the learning of text embeddings from overfitting to the current task, resulting in forgetting previous knowledge. When updating text embeddings, the text encoder $\mathcal{T}$ is kept frozen and the structured semantic prompts are tuned for knowledge adaptation.
III-C2 Project-Then-Learn
@@
To transfer the knowledge from text to image space effectively, we use the text embeddings to guide the learning of $\mathcal{G}_{t}$ by projecting image features on the text space and then updating the ReID model, i.e., the image encoder, to accumulate new knowledge. Formally,
$\mathcal{L}_{id}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp\big({f^{(t)}_{i}}^{\top}\cdot{{f^{S}_{j}}^{(t)}}\big)}{\sum\limits_{k=1}^{N^{(t)}}\exp\big({f^{(t)}_{i}}^{\top}\cdot{{f^{S}_{k}}^{(t)}}\big)},$ (4)
$q_{j}=\left\{\begin{aligned} &1-\epsilon+\frac{\epsilon}{N^{(t)}}&&,\ j=y_{i}^{(t)}\\ &\frac{\epsilon}{N^{(t)}}&&,\ {\rm otherwise}\end{aligned}\right$., (5)
@@
where $f_{i}^{(t)}$ means the image feature of the $i$-th image with the identity label $y_{i}^{(t)}$, $i\in[1,n^{(t)}]$, and $n^{(t)}$ is the total number of the images in $\mathcal{D}^{(t)}_{train}$. ${f^{S}}^{(t)}$ is the adapted text embeddings that is initialized with the L2 normalized text embeddings $\mathbf{T}^{(t)}$, and ${f_{j}^{S}}^{(t)}$ denotes the adapted text embedding of the $j$-th identity. $\epsilon$ is a small constant for label smoothing regularization [40], which is simply set to 0.1. Notably, Eq. 4 is effective in fine-tuning the image encoder $\mathcal{G}_{t}$ together with text embeddings. As a result, with the help of the slow-paced learner, both image encoder and text embeddings can be slowly updated so that new knowledge is accumulated without forgetting the old.
Since the $\mathcal{L}_{id}$ loss uses the adapted text embeddings, the learned text semantics from the first stage can be gradually lost. Therefore, we also use the $\mathcal{L}_{proj}$ loss [12], which keeps the learned text embeddings frozen and uses them to guide the learning of the image encoder. $\mathcal{L}_{proj}$ can alleviate the loss of text semantics in the knowledge adaptation and projection process. It is formulated as:
$\mathcal{L}_{proj}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp\big({f^{(t)}_{i}}^{\top}\cdot{{f^{T}_{j}}^{(t)}}\big)}{\sum\limits_{k=1}^{N^{(t)}}\exp\big({f^{(t)}_{i}}^{\top}\cdot{{f^{T}_{k}}^{(t)}}\big)},$ (6)
@@
III-C3 Alternative Variant
Benefiting from CLIP and our proposed SSP, text and image semantics are aligned in the shared latent space. Intuitively, image representations that are well aligned with text embeddings can also be used in Eq. 4 to guide knowledge transfer and accumulation. Specifically, we can average all the extracted image features of each identity to obtain image representations $\mathbf{V}^{(t)}$ at the beginning of the second stage training. Formally,
$\mathbf{V}^{(t)}={\left\{{f^{V}_{j}}^{(t)}=\frac{1}{\left|X^{(t)}_{j}\right|}\sum\limits_{f^{(t)}\in X^{(t)}_{j}}f^{(t)}\ \Bigg|\ j\in\left[1,N^{(t)}\right]\right\}},$ (7)
where $X^{(t)}_{j}$ denotes the set of all person images with the same identity label $j$, and |||*| denotes the cardinal number of the set. Then, L2 normalized $\mathbf{V}^{(t)}$ is used for the initialization of the classifier in Eq. 4. In Sec. IV-D, we thoroughly investigate the efficacy of our designs, including knowledge adaptation with text embeddings (w/ KA-T), text-aligned image representations (w/ KA-I), random initialization (by default), and slow-paced learner (w/ SL).
III-D Details of Training and Inference
Overall, two stages are involved in each step $t$ during the lifelong evolution of the model. In the first stage, both image encoder $\mathcal{G}_{t-1}$ and text encoder $\mathcal{T}$ are frozen, and Eq. 2 is adopted for the optimization of SSP. In the second stage, $\mathcal{G}_{t}$ is optimized with the guidance of the learned text embeddings. The losses involved in the second stage are formulated as:
$\mathcal{L}_{stage2}=\lambda_{1}\mathcal{L}_{proj}+\lambda_{2}\left(\mathcal{L}^{b}_{id}+\mathcal{L}_{id}\right)+\lambda_{3}\left(\mathcal{L}^{b}_{tri}+\mathcal{L}_{tri}\right),$ (8)
wherein $\mathcal{L}^{b}_{id}$ and $\mathcal{L}^{b}_{tri}$ denote the identity classification loss [14] and triplet loss [41] applied on the image features before the last projecting layer of the CLIP image encoder. $\mathcal{L}_{id}$ and $\mathcal{L}_{tri}$ are applied to the representations after projection. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are coefficients to balance between different losses. After training in step $t$, only the image encoder $\mathcal{G}_{t}$ is used for ReID evaluation. The cosine distances between two person images as measured during inference.
IV Experiments
IV-A Datasets and Evaluation Protocol
IV-A1 Seen Datasets
Two widely-used conventional same-cloth datasets, i.e., Market-1501 [42] and MSMT17 [43], and two widely-used cloth-changing datasets, i.e., LTCC [6] and PRCC [8], are used in our proposed LReID-Hybrid benchmarks for training. The dataset statistics are shown in Tab. I. We design six training orders to imitate different situations comprehensively as follows,